Opened 4 years ago

Last modified 3 years ago

#504 new flight processing

THL13/01, flight day 281/2013, Thales

Reported by: knpa Owned by: edfi
Priority: immediate Milestone: 2013 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/uk/THL13_01-2013_281_Thales

Data arrived from ARSF via network transfer on 10/10/2013

THL13/01 Scientific objective: Commercial flight to test Thales instrument

THL13/01 Priority: urgent

PI: Robert Booker

Sensors:

Camera (Copied to FTP 16/10/2013)
Eagle (Copied to FTP 16/10/2013)
Hawk (Copied to FTP 16/10/2013)
LiDAR (Dispatched 30/10/2013)

Change History (18)

comment:1 Changed 4 years ago by tipo

Processed nav, no problems found.
Processed Hyperspectral, has a bunch of exxtra bands containing no data, so alternate bands need to be used when generating tiffs. Lines 1-5 are test lines.

SCTs:

Line Hawk Eagle
06 -0.01 -0.07
07 -0.01 -0.07
08 -0.01 -0.07
Last edited 4 years ago by tipo (previous) (diff)

comment:2 Changed 4 years ago by tipo

Created hyperspectral delivery, but it's missing the logsheet as we don't have one yet.

Last edited 4 years ago by tipo (previous) (diff)

comment:3 Changed 4 years ago by tipo

Added logsheet.

comment:4 Changed 4 years ago by edfi

D/C complete one missing screen shot which has not been added. Mapped files are zipping

comment:5 Changed 4 years ago by tipo

Added missing screenshot to hyperspectral delivery.
reprocessed RCD using new calibration and created new delivery.

comment:6 Changed 4 years ago by edfi

Camera check complete, ready to deliver

comment:7 Changed 4 years ago by tipo

Deliveries copied to ftp ARSF15 (16/10/13)

comment:8 Changed 4 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

comment:9 Changed 4 years ago by tipo

Processed LiDAR. Line 3b has zero returns, otherwise it all seems good.

comment:10 Changed 3 years ago by stgo

D/Cd LiDAR, line 2 has an unclassified cloud at the beginning. Once this has been rectified the DEM and readme will need to be recreated.

comment:11 Changed 3 years ago by tipo

Fixed classification and remade delivery.

comment:12 Changed 3 years ago by stgo

Rechecked LiDAR, ready to deliver.

comment:13 Changed 3 years ago by tipo

LiDAR dispatched 30/10/2013

comment:14 Changed 3 years ago by tipo

  • Description modified (diff)

comment:15 Changed 3 years ago by emca

Note: the LiDAR should have been processed using the more recent calibration (from 04/11/2013).

Reprocessing this data using the new calibration information.

comment:16 Changed 3 years ago by emca

Recreated the LAS files, and copied the classification over from the original delivery.

Delivery ready to be checked.

comment:17 Changed 3 years ago by tipo

Checked LiDAR. No problems found.

comment:18 Changed 3 years ago by emca

Reprocessed LiDAR Data made available to Robert Booker via ftp

Note: See TracTickets for help on using tickets.